A family is on a road trip and there average speed is 60mi/h they travel a total distance of 500 miles, how long did it take them to get to their destination?

Answer by TimothyLamb(4379)

---
s = d / t
t = d / s
t = 500/60
t = 8.333 hours = 8 hours and 20 minutes
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php